# FedADMM-InSa: An Inexact and Self-Adaptive ADMM for Federated Learning

### v1

@@
Abstract. Federated learning (FL) is a promising framework for learning from distributed data while maintaining privacy. The development of efficient FL algorithms encounters various challenges, including heterogeneous data and systems, limited communication capacities, and constrained local computational resources. Recently developed FedADMM methods show great resilience to both data and system heterogeneity. However, they still suffer from performance deterioration if the hyperparameters are not carefully tuned. To address this issue, we propose an inexact and self-adaptive FedADMM algorithm, termed FedADMM-InSa. First, we design an inexactness criterion for the clients’ local updates to eliminate the need for empirically setting the local training accuracy. This inexactness criterion can be assessed by each client independently based on its unique condition, thereby reducing the local computational cost and mitigating the undesirable straggle effect. The convergence of the resulting inexact ADMM is proved under the assumption of strongly convex loss functions. Additionally, we present a self-adaptive scheme that dynamically adjusts each client’s penalty parameter, enhancing algorithm robustness by mitigating the need for empirical penalty parameter choices for each client. Extensive numerical experiments on both synthetic and real-world datasets are conducted. As validated by some numerical tests, our proposed algorithm can reduce the clients’ local computational load significantly and also accelerate the learning process compared to the vanilla FedADMM.
@@ 1 Introduction
With the increasing volume of data generated by massively distributed devices and organizations, traditional centralized deep learning paradigms encounter challenges in terms of data collection, privacy concerns, and scalability. To address the above challenges, federated learning (FL) [21, 22, 26] has emerged as a promising paradigm and has gained significant attention in recent years. In FL, multiple clients (devices) in a distributed environment collaboratively train a neural network model under the coordination of a central server, without centralizing their data. The training process of FL consists of several communication rounds. For example, in the most commonly used FedAvg [26] algorithm, at every communication round, each client computes a model update on its local data using the latest copy of the global model parameters and then sends the model update to the central server. The server aggregates these updates by averaging to update the global model parameters, which are then sent to all clients. Despite the fact that no raw data is transmitted, there is still a risk of privacy breaches, such as data reconstruction attacks [7, 34, 35, 37]. One of the key properties of FL is that the clients are massively distributed and have limited communication capacity [18, 26, 40]. To decrease the number of communication rounds needed to train a model, FedAvg suggests increasing local computation on each client. Unlike traditional distributed optimization [43] where consensus is performed after every local gradient computation, in FedAvg, the clients perform multiple full-batch or stochastic gradient descent steps before these local model parameters are aggregated in order to update the global model parameters. However, simply increasing local gradient descent steps as in FedAvg would cause the client drift [19] issue due to the inherent data and system heterogeneity in FL [18, 26]. The client drift issue comes from the fact that clients in FL normally possess unbalanced and non-independent and identically distributed (non-IID) local datasets, which leads to inconsistent minima of the local and global objective functions [19, 32]. As a result, the aggregated global model suffers from deviation and does not converge to its true optimum [24]. To address this, FedProx [23] was proposed by adding an extra proximal term to the client’s loss function to constrain the client’s local model to be close to the server’s global model. However, the extra proximal term might degrade the training performance unless carefully tuned [31]. In [19], a method called Scaffold is proposed to reduce client drift by using variance reduction to correct each client’s local updates. However, this method doubles the size of variables that need to be communicated between the server and clients, which is not ideal given the limited communication resources in FL. Recent findings show that FedADMM methods [1, 13, 33, 41, 42] based on the alternating direction method of multipliers (ADMM) [4, 9] are inherently resilient to heterogeneous clients in FL. FedADMM leverages dual variables to tackle statistical heterogeneity and accommodates system heterogeneity by tolerating variable amounts of work performed by clients. FedADMM maintains identical communication costs per round as FedAvg/Prox and generalizes them using the augmented Lagrangian with dual variables and an extra quadratic penalty term as the client’s local loss function. This modification strikes a balance between updating the client’s local model and staying consistent with the global model [13]. Despite the above-mentioned nice features, current FedADMM methods still suffer from performance deterioration and the straggler effect [18, 30] if the hyperparameters are not carefully tuned, especially the amount of local training workload and the choice of the penalty parameter. In particular, FedADMM methods require clients to perform a certain amount of local training workload, which corresponds to solving the ADMM subproblem inexactly. This is implemented empirically by either commanding clients to solve the subproblem to a constant given accuracy [13, 41, 42] or performing a fixed number of (stochastic) gradient descent steps [1, 33]. However, the accuracy of this solution significantly impacts the effectiveness of the algorithm [11]. Meanwhile, empirically assigning the same amount of local training workload overlooks the heterogeneity in clients’ data and systems (e.g., non-IID datasets and varying computational resources). This oversight may also cause a severe straggler effect, as waiting for the resource-constrained clients to finish the overload work will slow down the training process, and simply dropping them will lead to a biased global model [3, 42]. Another critical issue involves the selection of the penalty parameter of the quadratic term in the augmented Lagrangian functions. It has been shown in ADMM applications [14, 28, 36] that the efficiency of ADMM heavily depends on the penalty parameter. If the penalty parameter is chosen too small or too large, the solution time can increase significantly. The problem is more complicated in FedADMM as each client can use a different penalty parameter, and an inappropriate choice of the penalty parameter can potentially deteriorate the performance of FedADMM methods. To address the aforementioned issues, we propose an inexact and self-adaptive FedADMM algorithm, referred to as FedADMM-InSa. Firstly, we introduce an easy-to-implement and adaptive inexactness criterion to guide the client’s local training. Our approach eliminates the need to manually set gradient descent steps or predefine a constant accuracy. This provides each client with the flexibility to solve its subproblem inexactly based on its unique situation, thereby eliminating the potential straggler effect. Furthermore, we design a scheme to dynamically adjust each client’s penalty parameter based on the discrepancy between its local model parameters and the global model parameters, avoiding unexpected performance deterioration due to improperly chosen fixed penalty parameters. Overall, our main contributions are as follows:
@@
    We propose an inexactness criterion that enables each client to dynamically adjust the precision of local training in each communication round. There is no need to empirically set the number of gradient descent steps or perceived constant accuracy a priori. This flexibility allows our algorithm to better adapt to the heterogeneous clients and datasets, save local computational resources, and mitigate the straggler effect.
@@
    The convergence of our proposed algorithm using the inexactness criterion is analyzed. Extensive numerical experiments demonstrate the improved performance of our proposed inexactness criterion and self-adaptive penalty adjusting scheme. As validated by some numerical tests, our proposed algorithm reduces the clients’ local computational load by 64.3% while accelerating the learning process when compared to the vanilla FedADMM.
@@ 2.2 ADMM
The ADMM is a variant of the classic augmented Lagrangian method (ALM) [17, 27], where the subproblem of each ALM iteration is decomposed into two parts and then solved in a Gauss-Seidel manner. The ADMM possesses a distinct advantage as the decomposed subproblems are typically much easier than those encountered in the ALM, allowing for the exploitation of inherent properties and structures within the model under investigation. Moreover, the ADMM usually exhibits satisfactory numerical performance without the need for specific initial iterates. These characteristics make the ADMM a benchmark algorithm across various domains, including image processing [15], statistical learning [12, 38], and optimal control problems [10, 11, 29, 39], etc. For a more comprehensive understanding of the ADMM, one can refer to [4, 8].
@@ 2.3 Vanilla FedADMM
In this subsection, we present the application of ADMM in the context of FL, leading to the derivation of the vanilla FedADMM [42]. Subsequently, we systematically analyze the challenges and limitations associated with each iterative step of the vanilla FedADMM. To apply the ADMM to address the FL problem (2.1), we first introduce auxiliary variables $u_{i}=z$, $i\in[\mathrm{m}]$, and rewrite problem (2.1) into the following consensus setting:
@@
It is clear that problem (2.5) is equivalent to (2.1) in the sense that their optimal solutions coincide. The formulation of problem (2.5) naturally fits into the FL setting, where $u_{i}\in\mathbb{R}^{n}$ can be interpreted as the local model parameters held by client $i\in[m]$, and $z\in\mathbb{R}^{n}$ as the global model parameters held by the server. Therefore, in the following discussions, we concentrate on solving the optimization problem (2.5) instead of (2.1), as they are equivalent. For problem (2.5), we define its augmented Lagrangian function as
$L_{\alpha,\beta}(u,\lambda,z)=\sum_{i=1}^{m}{\alpha_{i}}L_{\beta_{i}}(u_{i},\lambda_{i},z),~{}\text{with}~{}L_{\beta_{i}}(u_{i},\lambda_{i},z)=f_{i}(u_{i})-\lambda_{i}^{\top}\left(u_{i}-z\right)+\frac{\beta_{i}}{2}\left\|u_{i}-z\right\|^{2}.$ (2.6)
@@
The $u$-subproblem (2.8a) and the $\lambda$-subproblem (2.8b) correspond to the client’s local update process and can be performed in parallel. The $\lambda$-subproblem (2.8b) is simple and has an analytical form. However, the $u$-subproblem (2.8a) is challenging due to the use of high-dimensional and nonlinear neural networks (such as the ResNets [16]) in FL. It generally lacks a closed-form solution. As a result, the subproblem (2.8a) should be solved iteratively and inexactly, and the implementation of the FedADMM (2.3) must be embedded by an internal iterative process for the subproblem (2.8a) In the vanilla FedADMM method, the clients normally use gradient-based methods for solving (2.8a) to a certain precision. For example, client $i$ can implement a prescribed number of $E$ steps of the full-batch gradient descent as follows:
$u_{i}^{k,e+1}=u_{i}^{k,e}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}\left(u_{i}^{k,e},\lambda_{i}^{k},z^{k}\right),\quad e=0,1,\ldots,E-1,$ (2.9)
@@
where $\eta_{i}>0$ is the learning rate (also known as the step size). Then, client $i$ sets $u_{i}^{k+1}\coloneqq u_{i}^{k,E}$ and uses it to update $\lambda_{i}^{k+1}$ by (2.8b). While other optimization schemes like stochastic gradient descent and quasi-Newton methods such as L-BFGS [25] are plausible, the accuracy of this inexact solution significantly influences the algorithm’s effectiveness [11]. A notable mathematical problem arises concerning the determination of an appropriate inexactness criterion for solving the subproblem (2.8a) inexactly. Current FedADMM methods often require an empirically perceived precision in advance. For instance, clients may be instructed to perform a fixed number of (stochastic) gradient descent steps [1, 33] or solve the subproblem up to a constant predefined accuracy [13, 41, 42]. However, empirically assigning the same amount of local training workload is not a good strategy for heterogeneous clients with non-IID data and varying computational resources. For instance, assigning too much workload may inefficiently utilize local computational resources and cause a severe straggler effect, especially among clients with limited computational resources. Moreover, there is no necessity to pursue excessively accurate solutions for (2.8a), particularly when the iterates are still far from the solution point. Additionally, current FedADMM methods employ a fixed penalty parameter $\beta$ throughout the training process, without considering the heterogeneity of clients in FL. If the penalty parameter is inappropriately chosen at the beginning, the method’s efficiency and performance would deteriorate without remedy. To tackle the above-mentioned issues, we design a readily implementable and appropriately accurate inexactness criterion for solving subproblem (2.8a) and a self-adaptive scheme to adjust each client’s penalty parameter $\beta_{i}$ in Section 3. As a result, an inexact and self-adaptive FedADMM is proposed, enabling clients to dynamically adjust local training precision each round, eliminating the need to preset accuracy. This flexibility enhances the adaptability of our algorithm to diverse clients and datasets, saves local computational resources, and mitigates the straggler effect. In addition, the self-adaptive penalty parameter scheme improves the robustness of our algorithm by eliminating the risk associated with inappropriate pre-selection of clients’ penalty parameters.
@@
The server can also incorporate the partial client participation strategy by selecting a subset of clients $\mathcal{M}^{k}\subseteq[m]$ to calculate updated parameters $u_{i}^{k+1}$ and $\lambda_{i}^{k+1}$ using (2.8a) and (2.8b) at the $(k+1)$-th communication round. Note from (2.12) that the clients can provide the server with calculated $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ instead of $u_{i}^{k+1}$ and $\lambda_{i}^{k+1}$, thereby reducing the communication cost. Meanwhile, for the unselected clients $i\notin\mathcal{M}^{k}$, the server can simply use their previously communicated local model parameters by setting $u_{i}^{k+1}\coloneqq u_{i}^{k}$ and $\lambda_{i}^{k+1}\coloneqq\lambda_{i}^{k}$. Finally, after receiving updated parameters $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and $\beta_{i}$ from the clients, the server can get the updated global model parameters $z^{k+1}$ by (2.12). With the above analysis of the client update and the server aggregation processes, we summarize the vanilla FedADMM algorithm based on (2.3) in Algorithm 1.
@@
6:        Set $u_{i}^{k,0}\coloneqq z^{k}$.
7:        for each local update step $e=0,1,\ldots,E-1$ do
8:           $u_{i}^{k,e+1}=u_{i}^{k,e}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(u_{i}^{k,e},\lambda_{i}^{k},z^{k})$.
@@
16:        Send $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}}$ to the server.
17:     end for
18:     Server side:
19:     Update $z^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
20:  end for

### v3

@@
Abstract
Federated learning (FL) is a promising framework for learning from distributed data while maintaining privacy. The development of efficient FL algorithms encounters various challenges, including heterogeneous data and systems, limited communication capacities, and constrained local computational resources. Recently developed FedADMM methods show great resilience to both data and system heterogeneity. However, they still suffer from performance deterioration if the hyperparameters are not carefully tuned. To address this issue, we propose an inexact and self-adaptive FedADMM algorithm, termed FedADMM-InSa. First, we design an inexactness criterion for the clients’ local updates to eliminate the need for empirically setting the local training accuracy. This inexactness criterion can be assessed by each client independently based on its unique condition, thereby reducing the local computational cost and mitigating the undesirable straggle effect. The convergence of the resulting inexact ADMM is proved under the assumption of strongly convex loss functions. Additionally, we present a self-adaptive scheme that dynamically adjusts each client’s penalty parameter, enhancing algorithm robustness by mitigating the need for empirical penalty parameter choices for each client. Extensive numerical experiments on both synthetic and real-world datasets have been conducted. As validated by some tests, our FedADMM-InSa algorithm improves model accuracy by 7.8% while reducing clients’ local workloads by 55.7% compared to benchmark algorithms.
keywords:
Federated learning , ADMM , inexactness criterion , client heterogeneity
[…]
[label1] organization=Chair for Dynamics, Control, Machine Learning and Numerics – Alexander von Humboldt Professorship, Department of Mathematics, Friedrich-Alexander-Universität Erlangen-Nürnberg, addressline=Cauerstrasse 11, city=Erlangen, postcode=91058, country=Germany
[…]
[label2] organization=Chair of Computational Mathematics, Fundación Deusto, addressline=Avenida de las Universidades, 24, city=Bilbao, postcode=48007, state=Basque Country, country=Spain
[…]
[label3] organization=Departamento de Matemáticas, Universidad Autónoma de Madrid, addressline=Ciudad Universitaria de Cantoblanco, city=Madrid, postcode=28049, country=Spain
@@ 1 Introduction
With the increasing volume of data generated by massively distributed devices and organizations, traditional centralized deep learning paradigms encounter challenges in terms of data collection, privacy concerns, and scalability. To address the above challenges, federated learning (FL) (Li et al., 2021, 2020a; McMahan et al., 2017) has emerged as a promising paradigm and has gained significant attention in recent years.
1.1 Background
In FL, multiple clients (devices) in a distributed environment collaboratively train a neural network model under the coordination of a central server, without centralizing their data. The training process of FL consists of several communication rounds. For example, in the most commonly used FedAvg (McMahan et al., 2017) algorithm, at every communication round, each client computes a model update on its local data using the latest copy of the global model parameters and then sends the model update to the central server. The server aggregates these updates by averaging to update the global model parameters, which are then sent to all clients. Despite the fact that no raw data is transmitted, there is still a risk of privacy breaches, such as data reconstruction attacks (Geng et al., 2023; Wang et al., 2023b; Xiao et al., 2024; Yang et al., 2023).
One of the key properties of FL is that the clients are massively distributed and have limited communication capacity (Kairouz et al., 2021; McMahan et al., 2017; Zhang et al., 2022). To decrease the number of communication rounds needed to train a model, FedAvg suggests increasing local computation on each client. Unlike traditional distributed optimization (Zinkevich et al., 2010) where consensus is performed after every local gradient computation, in FedAvg, the clients perform multiple epochs using full-batch or stochastic gradient descents before these local model parameters are aggregated in order to update the global model parameters.
1.2 Current Challenges
Nevertheless, simply increasing local training epochs as in FedAvg would cause the client drift (Karimireddy et al., 2020) issue due to the inherent data and system heterogeneity in FL (Kairouz et al., 2021; McMahan et al., 2017). The client drift issue comes from the fact that clients in FL normally possess unbalanced and non-independent and identically distributed (non-IID) local datasets, which leads to inconsistent minima of the local and global objective functions (Karimireddy et al., 2020; Wang et al., 2020). As a result, the aggregated global model suffers from deviation and does not converge to its true optimum (Li et al., 2019). To address this, FedProx (Li et al., 2020b) was proposed by adding an extra proximal term to the client’s loss function to constrain the client’s local model to be close to the server’s global model. However, the extra proximal term might degrade the training performance unless carefully tuned (Wang et al., 2022). In (Karimireddy et al., 2020), a method called Scaffold is proposed to reduce client drift by using variance reduction to correct each client’s local updates. However, this method doubles the size of variables that need to be communicated between the server and clients, which is not ideal given the limited communication resources in FL.
Recent findings show that FedADMM methods (Acar et al., 2021; Gong et al., 2022; Wang et al., 2023a; Zhang et al., 2021; Zhou and Li, 2023) based on the alternating direction method of multipliers (ADMM) (Boyd et al., 2011; Glowinski and Marroco, 1975) are inherently resilient to heterogeneous clients in FL. FedADMM leverages dual variables to tackle statistical heterogeneity and accommodates system heterogeneity by tolerating variable amounts of work performed by clients. FedADMM maintains identical communication costs per round as FedAvg/Prox and generalizes them using the augmented Lagrangian with dual variables and an extra quadratic penalty term as the client’s local loss function. This modification strikes a balance between updating the client’s local model and staying consistent with the global model.
1.3 Research Motivations
Despite the above-mentioned nice features, current FedADMM methods still suffer from performance deterioration and the straggler effect (Kairouz et al., 2021; Tan et al., 2023) if the hyperparameters are not carefully tuned, especially the amount of local training workload and the choice of the penalty parameter. In particular, FedADMM methods require clients to perform a certain amount of local training workload, which corresponds to solving the ADMM subproblem inexactly. This is implemented empirically by either commanding clients to solve the subproblem to a constant given accuracy (Gong et al., 2022; Zhang et al., 2021; Zhou and Li, 2023) or performing a fixed number of local epochs (Acar et al., 2021; Wang et al., 2023a). However, the accuracy of this solution significantly impacts the effectiveness of the algorithm (Glowinski et al., 2022). Meanwhile, empirically assigning the same amount of local training workload overlooks the heterogeneity in clients’ data and systems (e.g., non-IID datasets and varying computational resources). This oversight may also cause a severe straggler effect, as waiting for the resource-constrained clients to finish the overload work will slow down the training process, and simply dropping them will lead to a biased global model (Bonawitz et al., 2019; Zhou and Li, 2023).
Another critical issue involves the selection of the penalty parameter of the quadratic term in the augmented Lagrangian functions. It has been shown in ADMM applications (He et al., 2000; Song et al., 2016; Xu et al., 2017) that the efficiency of ADMM heavily depends on the penalty parameter. If the penalty parameter is chosen too small or too large, the solution time can increase significantly. The problem is more complicated in FedADMM as each client can use a different penalty parameter, and an inappropriate choice of the penalty parameter can potentially deteriorate the performance of FedADMM methods.
1.4 Main Contributions
To address the aforementioned issues, we propose an inexact and self-adaptive FedADMM algorithm, referred to as FedADMM-InSa. Firstly, we introduce an easy-to-implement and adaptive inexactness criterion to guide the client’s local training. Our approach eliminates the need to manually set local epochs or predefine a constant accuracy. This provides each client with the flexibility to solve its subproblem inexactly based on its unique situation, thereby eliminating the potential straggler effect. Furthermore, we design a scheme to dynamically adjust each client’s penalty parameter based on the discrepancy between its local model parameters and the global model parameters, avoiding unexpected performance deterioration due to improperly chosen fixed penalty parameters.
Overall, our main contributions are as follows:
@@ 1.4 Main Contributions
    We propose an inexactness criterion that enables each client to dynamically adjust the precision of local training in each communication round. There is no need to empirically set the number of local training epochs or perceived constant accuracy a priori. This flexibility allows our algorithm to better adapt to the heterogeneous clients and datasets, save local computational resources, and mitigate the straggler effect.
@@
    The convergence of our proposed algorithm using the inexactness criterion is analyzed. Extensive numerical experiments demonstrate the improved performance of our proposed inexactness criterion and self-adaptive penalty adjusting scheme. As validated by some numerical tests, our proposed algorithm reduces the clients’ local computational load by 55.7% while accelerating the learning process when compared to the vanilla FedADMM.
1.5 Organization
@@ 2.2 ADMM
Then, starting with an initial point $\{y^{0},\lambda^{0}\}$, for $k\geq 0$, the ADMM (Glowinski and Marroco, 1975) iteratively updates the variables $\left\{x,y,\lambda\right\}$ in the following way:
@@
The ADMM is a variant of the classic augmented Lagrangian method (ALM) (Hestenes, 1969; Powell, 1969), where the subproblem of each ALM iteration is decomposed into two parts and then solved in a Gauss-Seidel manner. The ADMM possesses a distinct advantage as the decomposed subproblems are typically much easier than those encountered in the ALM, allowing for the exploitation of inherent properties and structures within the model under investigation. Moreover, the ADMM usually exhibits satisfactory numerical performance without the need for specific initial iterates. These characteristics make the ADMM a benchmark algorithm across various domains, including image processing (He and Yuan, 2018), statistical learning (Goldstein et al., 2015; Yue et al., 2018), and optimal control problems (Glowinski et al., 2020, 2022; Song et al., 2023; Zhang et al., 2017), etc. For a more comprehensive understanding of the ADMM, one can refer to (Boyd et al., 2011; Glowinski, 2014).
@@ 2.3 Vanilla FedADMM
In this subsection, we present the application of ADMM in the context of FL, leading to the derivation of the vanilla FedADMM (Zhou and Li, 2023). Subsequently, we systematically analyze the challenges and limitations associated with each iterative step of the vanilla FedADMM.
To apply the ADMM to address the FL problem (2.1), we first introduce auxiliary variables $u_{i}=z$, $i\in[\mathrm{m}]$, and rewrite problem (2.1) into the following consensus setting:
@@
It is clear that problem (2.5) is equivalent to (2.1) in the sense that their optimal solutions coincide. The formulation of problem (2.5) naturally fits into the FL setting, where $u_{i}\in\mathbb{R}^{n}$ can be interpreted as the local model parameters held by client $i\in[m]$, and $z\in\mathbb{R}^{n}$ as the global model parameters held by the server. Therefore, in the following discussions, we concentrate on solving the optimization problem (2.5) instead of (2.1), as they are equivalent.
For problem (2.5), we define its augmented Lagrangian function as
$L_{\alpha,\beta}(u,\lambda,z)$ $=\sum_{i=1}^{m}{\alpha_{i}}L_{\beta_{i}}(u_{i},\lambda_{i},z),\text{ with}$ (2.6)
$L_{\beta_{i}}(u_{i},\lambda_{i},z)$ $=f_{i}(u_{i})-\lambda_{i}^{\top}\left(u_{i}-z\right)+\frac{\beta_{i}}{2}\left\|u_{i}-z\right\|^{2}.$
@@
The $u$-subproblem (2.8a) and the $\lambda$-subproblem (2.8b) correspond to the client’s local update process and can be performed in parallel. The $\lambda$-subproblem (2.8b) is simple and has an analytical form. However, the $u$-subproblem (2.8a) is challenging due to the use of high-dimensional and nonlinear neural networks (such as the ResNets (He et al., 2016)) in FL. It generally lacks a closed-form solution. As a result, the subproblem (2.8a) should be solved iteratively and inexactly, and the implementation of the FedADMM (2.8) must be embedded by an internal iterative process for the subproblem (2.8a)
In the vanilla FedADMM method, the clients normally use gradient-based methods for solving (2.8a) to a certain precision. For example, client $i$ first sets $\hat{u}_{i}^{k}\coloneqq z^{k}$, then implements a prescribed number of local epochs consisting of multiple gradient descent steps:
$\hat{u}_{i}^{k}\coloneqq\hat{u}_{i}^{k}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(\hat{u}_{i}^{k},\lambda_{i}^{k},z^{k}),$ (2.9)
where $\eta_{i}>0$ is the learning rate (also known as the step size). Then, client $i$ sets $u_{i}^{k+1}\coloneqq\hat{u}_{i}^{k}$ and uses it to update $\lambda_{i}^{k+1}$ by (2.8b). While other optimization schemes like stochastic gradient descent and quasi-Newton methods such as L-BFGS (Liu and Nocedal, 1989) are plausible, the accuracy of this inexact solution significantly influences the algorithm’s effectiveness (Glowinski et al., 2022). A notable mathematical problem arises concerning the determination of an appropriate inexactness criterion for solving the subproblem (2.8a) inexactly.
@@
Current FedADMM methods often require an empirically perceived precision in advance. For instance, clients may be instructed to perform a fixed number of training epochs (Acar et al., 2021; Wang et al., 2023a) or solve the subproblem up to a constant predefined accuracy (Gong et al., 2022; Zhang et al., 2021; Zhou and Li, 2023). However, empirically assigning the same amount of local training workload is not a good strategy for heterogeneous clients with non-IID data and varying computational resources. For instance, assigning too much workload may inefficiently utilize local computational resources and cause a severe straggler effect, especially among clients with limited computational resources. Moreover, there is no necessity to pursue excessively accurate solutions for (2.8a), particularly when the iterates are still far from the solution point. Additionally, current FedADMM methods employ a fixed penalty parameter $\beta$ throughout the training process, without considering the heterogeneity of clients in FL. If the penalty parameter is inappropriately chosen at the beginning, the method’s efficiency and performance would deteriorate without remedy.
To tackle the above-mentioned issues, we design a readily implementable and appropriately accurate inexactness criterion for solving subproblem (2.8a) and a self-adaptive scheme to adjust each client’s penalty parameter $\beta_{i}$ in Section 3. As a result, an inexact and self-adaptive FedADMM is proposed, enabling clients to dynamically adjust local training precision each round, eliminating the need to preset accuracy. This flexibility enhances the adaptability of our algorithm to diverse clients and datasets, saves local computational resources, and mitigates the straggler effect. In addition, the self-adaptive penalty parameter scheme improves the robustness of our algorithm by eliminating the risk associated with inappropriate pre-selection of clients’ penalty parameters.
@@
The server can also incorporate the partial client participation strategy by selecting a subset of clients $\mathcal{M}^{k}\subseteq[m]$ to calculate updated parameters $u_{i}^{k+1}$ and $\lambda_{i}^{k+1}$ using (2.8a) and (2.8b) at the $(k+1)$-th communication round. Note from (2.12) that the clients can provide the server with calculated $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ instead of $u_{i}^{k+1}$ and $\lambda_{i}^{k+1}$, thereby reducing the communication cost. Meanwhile, for the unselected clients $i\notin\mathcal{M}^{k}$, the server can simply use their previously communicated local model parameters by setting $u_{i}^{k+1}\coloneqq u_{i}^{k}$ and $\lambda_{i}^{k+1}\coloneqq\lambda_{i}^{k}$. Finally, after receiving updated parameters $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and $\beta_{i}$ from the clients, the server can get the updated global model parameters $z^{k+1}$ by (2.12).
With the above analysis of the client update and the server aggregation processes, we summarize the vanilla FedADMM algorithm based on (2.8) in Algorithm 1.
@@
6:        Set $\hat{u}_{i}^{k}\coloneqq z^{k}$.
7:        for each epoch $e=1,\ldots,E$ do
8:           Update $\hat{u}_{i}^{k}\coloneqq\hat{u}_{i}^{k}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(\hat{u}_{i}^{k},\lambda_{i}^{k},z^{k})$ for one epoch.
@@
16:     end for
17:     Server side:
18:     Update $z^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
19:  end for